# Ontological differentiation as a measure of semantic accuracy

### v2

@@ Abstract
Understanding semantic relationships within complex networks derived from lexical resources is fundamental for network science and language modeling. While network embedding methods capture contextual similarity, quantifying semantic distance based directly on explicit definitional structure remains challenging. Accurate measures of semantic similarity allow for navigation on lexical networks based on maximizing semantic similarity in each navigation jump (Semantic Navigation, SN). This work introduces Ontological Differentiation (OD), a formal method for measuring divergence between concepts by analyzing overlap during recursive definition expansion. The methodology is applied to networks extracted from the Simple English Wiktionary, comparing OD scores with other measures of semantic similarity proposed in the literature (cosine similarity based on random-walk network exploration). We find weak correlations between direct pairwise OD scores and cosine similarities across $\sim$ 2 million word pairs, sampled from a pool representing over 50% of the entries in the Wiktionary lexicon. This establishes OD as a largely independent, definition-based semantic metric, whose orthogonality to cosine similarity becomes more pronounced when low-semantic-content terms were removed from the dataset. Additionally, we use cumulative OD scores to evaluate paths generated by vector-based SN and structurally optimal Shortest Paths (SP) across networks. We find SN paths consistently exhibit significantly lower cumulative OD scores than shortest paths, suggesting that SN produces trajectories more coherent with the dictionary’s definitional structure, as measured by OD. Ontological Differentiation thus provides a novel, definition-grounded tool for analyzing semantic structure and validating navigation processes in lexical networks.
@@ I Introduction
In this paper, we introduce and formalize Ontological Differentiation (OD), a method designed to quantify, based on first principles, the divergence between concepts based directly on their recursive definition structure within a lexical network. The core idea is that the semantic distance between two concepts relates to how quickly and extensively their definitional expansions overlap. We operationalize this by iteratively “reading” the definitions of a starting set of concepts and applying specific cancelation rules when elements are repeated across different expansion branches. The process continues until a termination condition is met, and the final OD score is calculated as a weighted sum reflecting the depth at which cancelations occurred. We focus primarily on Strong Ontological Differentiation (SOD), where cancelations are triggered only by overlaps between elements originating from different starting concepts.
@@
Our primary goals are fourfold: (1) to provide a rigorous theoretical framework for defining OD as a meaningful measure of semantic proximity between terms; (2) to investigate the relationship between the definition-based SOD metric and the network-embedding-based cosine similarity derived from RIM; (3) to assess the robustness of these findings by applying our methods to networks constructed from the Simple English Wiktionary under three different definition processing schemes (ground-filtered, random removal, and targeted removal), simulating variations in corpus quality and authorship; and (4) having established the nature of the relationship between SOD and cosine similarity across these conditions, to apply cumulative SOD scores as an independent benchmark to evaluate the semantic coherence of paths generated by different navigation algorithms. Specifically, we use this framework to compare paths generated by Semantic Navigation (SN) [17], which optimizes RIM vector-based similarities, and standard Shortest Path (SP) navigation, which minimizes the overall path weight, using a large-scale sample of word pairs. By first understanding how SOD relates to cosine similarity, we can better interpret the results of using SOD to compare SN and SP paths.
@@
Our results demonstrate that, while the precise correlation between SOD and RIM cosine similarity varies with the definition processing scheme (ranging from weak negative in less processed networks to negligible in the targeted removal network), SOD consistently acts as a largely independent measure of definitional content. Furthermore, and critically for its utility as an evaluation metric, we find that SN paths consistently exhibit lower cumulative SOD scores than SP paths across all tested network conditions. This suggests that SOD quantitatively validates the intuition that SN generates more definitionally coherent trajectories through the semantic network, a finding robust to significant variations in the underlying definitional structure.
@@
The remainder of this paper is structured as follows: Section II details the theoretical framework of Ontological Differentiation and its variants. Section III describes the Semantic Navigation and Shortest Path algorithms used for comparison. Section IV outlines the three definition processing schemes leading to distinct datasets, network construction, vector computation, the SOD evaluation framework, the design of our large-scale experiment (including details on lexicon size and sampling), and an illustrative example. Section V first presents the comparative analysis of the relationship between direct SOD scores and RIM vector cosine similarity across the three conditions, and subsequently shows the results comparing the cumulative SOD scores of SN and SP paths. Finally, Section VI discusses the implications of our findings. Conceptual examples illustrating OD computation, together with detailed distributional plots for each dataset condition are provided in the Supplemental Material. In addition, the Supplemental Material includes a crucial extended analysis that investigates the confounding effects of network topology and path length, confirming the robustness of our central conclusions.
@@ II First-principle definition of semantic similarity
Let $\mathbf{U}$ be a set of elements $\{P_{1},...,P_{n}\}$ of size $n\in\mathbb{N}$, for which every element $P_{i}\in\mathbf{U}$ is itself a set of size $m_{i}\in\mathbb{N}$ (can vary per element) of elements in $\mathbf{U}$, that is $P_{i}=\{P_{i_{1}},...,P_{i_{m_{i}}}\}$ where every $P_{i_{j}}\in\mathbf{U}$. This represents a collection of concepts where each concept is defined by a set of other concepts in the collection.
@@ II.1 Ontological differentiation
The process of applying a specific cancelation rule $\kappa$ and termination rule $\tau$ to an initial set $\mathcal{A}$ is denoted $F_{\kappa\tau}(\mathcal{A})$. This process determines which elements get canceled at each level $n$ (from 00 to $\omega$) and the final expansion depth $\omega$. As a result, $F_{\kappa\tau}(\mathcal{A})$ yields a finite, recursive expansion of the terms included initially included in $\mathcal{A}$. We name the process $F_{\kappa\tau}(\mathcal{A})$ as the Ontological Differentiation of set $\mathcal{A}$.
@@ II.2 Types of ontological differentiation
The general OD framework allows for various types, each defined by specific cancelation and termination rules. Here we introduce Strong Ontological Differentiation (SOD) first, due to its intuitive alignment with how semantic divergence might be perceived based on definitional content, including an illustrative example of its calculation based on dictionary terms. For that reason, in this work we focus primarily on strong ontological differentiation (SOD). However, in this subsection we briefly describe a second type of OD, named Weak Ontological Differentiation (WOD), to illustrate the flexibility and generality of our OD framework, as well as Great Ontological Differentiation (GOD), which is key in bounding the number of recursion steps in the expansion process. Conceptual examples further illustrating the computation of WOD and GOD scores, based on abstract concepts, are provided in the Supplemental Material.
@@ II.2.1 Strong ontological differentiation (SOD)
We illustrate how SOD operates with the aforementioned rules by calculating the score for the pair of words ‘job’ and ‘business’. We will demonstrate the SOD calculation assuming that the definitions for these terms and their constituents are as provided in the “Definitions” column of Table II.2.1, as might be extracted from a lexical resource after a filtering process similar to that described later in Section IV.1.
The SOD process, detailed in Table II.2.1, unfolds by iteratively expanding these definitions. After each level of expansion, the SOD cancelation rule ($\kappa_{S}$) is applied by looking for elements appearing on both the ‘job’ and ‘business’ sides across all currently expanded levels. Subsequently, the SOD termination rule ($\tau_{S}$) is checked. In Table II.2.1, elements that are ultimately canceled by this process are shown in bold.
@@
      Expansion: Elements from $R^{1}(\text{job})$ and $R^{1}(\text{business})$ are expanded according to the corresponding definitions (see Table II.2.1).
@@
[…]
@@ II.2.2 Weak ontological differentiation (WOD)
Weak Ontological Differentiation (WOD) offers a less restrictive approach to measuring divergence. Unlike SOD, which only considers cross-side repetitions for cancelation, WOD cancels an element if it has appeared anywhere before in the overall expansion $R^{\omega}(\mathcal{A})$, regardless of its origin side or level. WOD uses the following cancelation rule, while its termination rule ($\tau_{W}$) is the same as that defined for SOD ($\tau_{S}$).
Definition II.10.
(WOD cancelation rule, $\kappa_{W}$) An instance of element $\alpha$ generated at level $n$ from $P_{s}\in\mathcal{A}$ is canceled if $\alpha$ has appeared at least once before in the expansion $R^{\omega}(\mathcal{A})$ up to the current overall level $n_{c}\leq\omega$ being considered, regardless of its origin ($P_{r}$) or its own generation level ($h\leq n_{c}$). Formally, if $\alpha\in R^{n}(P_{s})$ (where $n\leq n_{c}$) and there exists $\beta\in R^{h}(P_{r})$ (where $h\leq n_{c}$) such that $\alpha=\beta$ for any $(h,r)\neq(n,s)$, then $\alpha$ is canceled.
This implies that, in WOD, any element appearing at least twice anywhere within the total expansion $R^{\omega}(\mathcal{A})$ will be marked as canceled (except for its first appearance). The process terminates, identically to SOD, when one side’s expansion at a particular level becomes entirely canceled. WOD is presented here primarily to illustrate that SOD is one specific instantiation within the broader OD framework; other OD types can be defined by varying the cancelation or termination rules.
A detailed, step-by-step illustration of WOD is provided in the Supplemental Material. This appendix example, like others provided for conceptual clarity, presents a more exhaustive, visual walkthrough of the iterative expansion and cancelation process. Such detailed tracking is excellent for understanding the core logic, but it is not the most computationally efficient representation for large-scale application. The actual calculations performed in this study utilize an optimized algorithm that tracks sets of uncanceled and canceled elements at each level for greater efficiency, as detailed and implemented in the code available in our repository [19].
@@ II.2.3 Great ontological differentiation (GOD)
The level $\omega_{GOD}$, illustrated with an abstract example in the Supplemental Material, signifies the point beyond which no new concepts are introduced into the collective definitional expansion. Therefore, it serves as a practical upper bound for SOD/WOD: if the primary termination rule has not been met by $\omega_{GOD}+1$, the OD process can be halted. This prevents infinite loops and ensures the differentiation process has explored all potentially new definitional content before terminating due to stability rather than complete cancelation on one side.
@@ III Navigation methods
In this section, we describe two distinct strategies for finding a route between a source node $s$ and a target node $t$ in a semantic network: (1) semantic navigation (SN), which leverages high‐order connectivity via vector embeddings, and (2) shortest‐path (SP) navigation, which finds cost‐minimal routes in the graph.
@@ III.1 Graph representation
The graph is constructed from the initial directed definition links. First, for each headword $i$ defined by a set of $m_{i}$ tokens, a directed edge is drawn from $i$ to each token $j$. The network is then symmetrized by adding a reverse edge $j\to i$ for every original edge $i\to j$. This choice is crucial for modeling the real, navigable space of a digital lexicon. For instance, if entry B is in the definition of A, a directed link $A\to B$ exists. However, A may not appear in the definition of B, meaning the reverse link $B\to A$ is absent. A user navigating from A to B can, however, always return to A (e.g., via a “back” button). To capture this fundamental bidirectional navigability, we enforce symmetry. This process reflects the principle that a definitional connection, even if unidirectional, implies a mutual semantic association that can be traversed in both directions. While this symmetric network forms the basis of our main results, we conduct a parallel analysis on the original (non-symmetrized) directed network in the Supplemental Material to validate our findings under stricter assumptions.
Each edge $(i,j)$ carries a non-negative, symmetric weight $w_{ij}=w_{ji}$. If the original directed link $i\to j$ existed because $j$ was one of the $m_{i}$ symbols in the definition of $i$, the weight is set to $w_{ij}=1/m_{i}$. The corresponding reverse edge $j\to i$ is assigned the same weight.
@@ IV.1 Data acquisition and definition processing
Initially, the definitions corresponding to our working lexicon underwent a common set of rigorous filtering and cleaning processes aimed at isolating core semantic relationships and standardizing the lexical base. For entries exhibiting polysemy, we adopted the simplifying convention of using only the first definition provided in the Wiktionary entry. This choice, while expedient for large-scale processing, acknowledges that not all facets of a word’s meaning may be captured. Entries’ definitions were then filtered based on their likely part-of-speech (only nouns, verbs, and adjectives retained) to focus on core lexical concepts. Furthermore, heuristics were applied to filter terms potentially representing inflected forms and to identify lemmas for definition tokens, aiming to build the network primarily from base forms. The text was then tokenized into constituent words. These tokens were further filtered: very short or numerical tokens were removed. Finally, to prevent potential trivial results in the subsequent OD calculations, any instance of a headword appearing within its own processed definition was removed.

### v3

@@ Abstract
Understanding semantic relationships within complex networks derived from lexical resources is fundamental for network science and language modeling. While network embedding methods capture contextual similarity, quantifying semantic distance based directly on explicit definitional structure remains challenging. Accurate measures of semantic similarity allow for navigation on lexical networks based on maximizing semantic similarity in each navigation jump (Semantic Navigation, SN). This work introduces Ontological Differentiation (OD), a formal method for measuring divergence between concepts by analyzing overlap during recursive definition expansion. The methodology is applied to networks extracted from the Simple English Wiktionary, comparing OD scores with other measures of semantic similarity proposed in the literature (cosine similarity based on random-walk network exploration). We find weak correlations between direct pairwise OD scores and cosine similarities across $\sim$ 2 million word pairs, sampled from a pool representing over 50% of the entries in the Wiktionary lexicon. This establishes OD as a largely independent, definition-based semantic metric, whose orthogonality to cosine similarity becomes more pronounced when low-semantic-content terms were removed from the dataset. Additionally, we use cumulative OD scores to evaluate paths generated by vector-based SN and structurally optimal Shortest Paths (SP) across networks. We find SN paths consistently exhibit significantly lower cumulative OD scores than shortest paths, suggesting that SN produces trajectories more coherent with the dictionary’s definitional structure, as measured by OD. Ontological Differentiation thus provides a novel, definition-grounded tool for analyzing, validating, and potentially constructing navigation processes in lexical networks.
@@ I Introduction
In this paper, we introduce and formalize Ontological Differentiation (OD), a method designed to quantify, from first principles, the divergence between concepts using the recursive structure of their definitions within a lexical network. At an intuitive level, OD works by repeatedly expanding each initial concept into the terms that appear in its definition, and then expanding those terms in turn, producing successive layers of related words. When a term is repeated during this process, a so called cancellation occurs, signaling a point of semantic overlap. The framework allows for different variants based on specific cancellation rules. In Strong Ontological Differentiation (SOD), the focus of this work, cancellations are triggered only by overlaps between elements originating from different starting concepts. In contrast, Weak Ontological Differentiation (WOD) cancels an element if it has appeared anywhere before, regardless of its origin. The process continues until a termination condition is met, at which point the final OD score is calculated as a weighted sum reflecting the depth at which cancellations occurred. We present WOD primarily to illustrate the framework’s flexibility [19], (see Sec. II for the precise rules and a worked SOD example).
@@
Our primary goals are fourfold: (1) to provide a rigorous theoretical framework for defining OD as a meaningful measure of semantic proximity between terms; (2) to investigate the relationship between the definition-based SOD metric and the network-embedding-based cosine similarity derived from RIM; (3) to assess the robustness of these findings by applying our methods to networks constructed from the Simple English Wiktionary under three different definition processing schemes (ground-filtered, random removal, and targeted removal), simulating variations in corpus quality and authorship; and (4) having established the nature of the relationship between SOD and cosine similarity across these conditions, to apply cumulative SOD scores as an independent benchmark to evaluate the semantic coherence of paths generated by different navigation algorithms. Specifically, we use this framework to compare paths generated by Semantic Navigation (SN) [17], which optimizes RIM vector-based similarities, and standard Shortest Path (SP) navigation, which minimizes the overall path weight, using a large-scale sample of word pairs.
@@
Our results demonstrate that, while the precise correlation between SOD and RIM cosine similarity varies with the definition processing scheme, SOD consistently acts as a largely independent measure of definitional content, a conclusion supported by both large-scale quantitative analysis and focused qualitative examples that reveal the distinct semantic logic of each metric. Furthermore, and critically for its utility as an evaluation metric, we find that SN paths consistently exhibit lower cumulative SOD scores than SP paths across all tested network conditions. This suggests that SOD quantitatively validates the intuition that SN generates more definitionally coherent trajectories through the semantic network, a finding robust to significant variations in the underlying definitional structure, and it motivates our proof-of-concept exploration of Ontological Navigation, which shows that SOD can serve as a principled navigation heuristic in its own right.
Several prior methods also exploit dictionary or ontology definitions to measure semantic similarity. That is the case of, e. g., Extended Gloss Overlap (EGO) [20], where two glosses are treated as flat pieces of text, and similarity is measured by the number and length of overlapping word sequences between them, so similarity is a function of the direct word overlap between glosses. As a consequence, EGO is fundamentally a non-recursive, single-step comparison. Conversely, simDEF [21] represents, in turn, each definition as a bag of words (possibly extended to neighbors in an ontology) and compares these profiles via co-occurrence statistics. Therefore, it converts definitions into a vector space representation and then computes similarity using a geometric metric like cosine similarity. EGO and simDEF compute, therefore, similarity as the degree of static overlap between glosses or definition sets.
Alternatively, recent advances in artificial intelligence, such as node2vec [22], Graph Neural Networks, and Large Language Models, provide powerful embedding-based approaches to semantic similarity. These methods produce high-dimensional vector representations trained on distributional or structural data.
In contrast to the previous approaches, OD provides an entirely symbolic and definition-grounded process-based measure with two key features. First, it is fundamentally recursive and depth-weighted: expansions unfold level by level, and the score explicitly reflects not only how many overlaps occur but also how quickly and how abundantly they appear. Second, it is governed by explicit termination conditions, halting only when one concept’s definitional structure has been fully accounted for by the other. In this sense, the OD score captures the cumulative effort required for one concept to semantically reach the other, providing a conceptually distinct, combinatorial notion of semantic divergence compared to overlap-based or vector-space methods. Third, in contrast to other artificial-intelligence approaches, OD requires no training and yields an interpretable semantic distance based on recursive definitional overlap. Thus, OD distinguishes as a complementary technique since it offers a transparent benchmark for semantic coherence that can inform, evaluate, or be integrated with embedding-based techniques.
@@
The remainder of this paper is structured as follows: Section II details the theoretical framework of Ontological Differentiation and its variants. Section III describes the Semantic Navigation and Shortest Path algorithms used for comparison. Section IV outlines the three definition processing schemes leading to distinct datasets, network construction, vector computation, the SOD evaluation framework, the design of our large-scale experiment (including details on lexicon size and sampling), and an illustrative example. Section V presents our main findings: it first provides a quantitative and qualitative comparison between SOD and cosine similarity, then shows the results comparing the cumulative SOD scores of SN and SP paths, and culminates in a proof-of-concept exploration of Ontological Navigation. Finally, Section VI discusses the implications of our findings. In addition, the SM [19] provides conceptual examples illustrating OD computation, detailed distributional plots, an analysis of the method’s computational scalability, a study on its sensitivity to polysemy, and a crucial extended analysis that investigates the confounding effects of network topology and path length.
@@ II First-principle definition of semantic similarity
Let $\mathbf{U}$ be a set of elements $\{P_{1},...,P_{n}\}$ of size $n\in\mathbb{N}$, for which every element $P_{i}\in\mathbf{U}$ is itself a set of size $m_{i}=|P_{i}|$ (can vary per element) of elements in $\mathbf{U}$, that is $P_{i}=\{P_{i_{1}},...,P_{i_{m_{i}}}\}$ where every $P_{i_{j}}\in\mathbf{U}$. This represents a collection of concepts where each concept is defined by a set of other concepts in the collection.
@@ II.1 Ontological differentiation
This definition is intentionally stated in general terms, abstracting away from the specifics of which expansions are compared. To see how such rules operate on expansions and how overlaps trigger cancellations under SOD and WOD, we refer the interested reader to the worked examples in Sec. II.2.1 and in the SM [19].
@@
The process of applying a specific cancelation rule $\kappa$ and termination rule $\tau$ to an initial set $\mathcal{A}$ is denoted $F_{\kappa\tau}(\mathcal{A})$. This process determines which elements get canceled at each level $n$ (from 0 to $\omega$) and the final expansion depth $\omega$. As a result, $F_{\kappa\tau}(\mathcal{A})$ yields a finite, recursive expansion of the terms initially included in $\mathcal{A}$. We name the process $F_{\kappa\tau}(\mathcal{A})$ as the Ontological Differentiation of set $\mathcal{A}$.
@@ II.2 Types of ontological differentiation
The general OD framework allows for various types, each defined by specific cancelation and termination rules. Here we introduce Strong Ontological Differentiation (SOD) first, due to its intuitive alignment with how semantic divergence might be perceived based on definitional content. The primary distinction between the variants we discuss lies in what constitutes a meaningful repetition. Consider the expansion originating from two initial concepts, $P_{A}$ and $P_{B}$. SOD employs a strict, cross-branch rule: an element is canceled only if it appears in the expansion of $P_{A}$ and also appears in the expansion of $P_{B}$. This directly measures the convergence of distinct definitional paths. For that reason, in this work we focus primarily on SOD.
However, to illustrate the flexibility of our framework, besides SOD, we also briefly describe a second type, Weak Ontological Differentiation (WOD), which uses a more general rule: an element is canceled if it has appeared anywhere before in the total expansion, regardless of which branch it originated from. Finally, we introduce Great Ontological Differentiation (GOD), which is key in bounding the number of recursion steps in the expansion process. Conceptual examples further illustrating the computation of WOD and GOD scores, based on abstract concepts, are provided in the SM [19].
@@ II.2.1 Strong ontological differentiation (SOD)
We illustrate how SOD operates with the aforementioned rules by calculating the score for the pair of words ‘job’ and ‘business’. We will demonstrate the SOD calculation assuming that the definitions for these terms and their constituents are as provided in the “Definitions” column of Table 1, as might be extracted from a lexical resource after a filtering process similar to that described later in Section IV.1.
The SOD process, detailed in Table 1, unfolds by iteratively expanding these definitions. After each level of expansion, the SOD cancelation rule ($\kappa_{S}$) is applied by looking for elements appearing on both the ‘job’ and ‘business’ sides across all currently expanded levels. Subsequently, the SOD termination rule ($\tau_{S}$) is checked. In Table 1, elements that are ultimately canceled by this process are shown in bold.
@@
      Expansion: Elements from $R^{1}(\text{job})$ and $R^{1}(\text{business})$ are expanded according to the corresponding definitions (see Table 1).
@@
      Termination check: After all cancelations are resolved, all elements of the multiset $R^{1}(\text{job})=\{\text{pay, money, service}\}$ are now marked as canceled. Therefore, the termination condition $\tau_{S}$ is met. The process stops, and the final score is calculated as shown in the table.
@@ II.2.2 Weak ontological differentiation (WOD)
For completeness, we note a less restrictive variant, Weak Ontological Differentiation (WOD). Unlike SOD, which cancels elements only when they appear in expansions of different initial concepts (cross-branch repetitions), WOD cancels any repeated element regardless of its origin branch or level. This variant illustrates that OD is a general framework with multiple possible cancellation rules.
A full definition and a detailed, step-by-step worked example of WOD (including its termination and scoring) are provided in the SM [19]. This appendix example, like others provided for conceptual clarity, presents a more exhaustive, visual walkthrough of the iterative expansion and cancelation process. Such detailed tracking is excellent for understanding the core logic but is not the computationally efficient method used for large-scale application. The actual calculations performed in this study utilize an optimized algorithm that tracks sets of uncanceled and canceled elements at each level. The implementation of this algorithm is available in our repository [23], and a detailed analysis of its computational cost, scalability and termination guarantees is presented in the SM [19].
@@ II.2.3 Great ontological differentiation (GOD)
The level $\omega_{GOD}$, illustrated with an abstract example in the SM [19], signifies the point beyond which no new concepts are introduced into the collective definitional expansion. Therefore, it serves as a practical upper bound for SOD/WOD: if the primary termination rule has not been met by $\omega_{GOD}+1$, the OD process can be halted. This prevents infinite loops and ensures the differentiation process has explored all potentially new definitional content before terminating due to stability rather than complete cancelation on one side.
@@ III Navigation methods
In this section, we describe two distinct strategies for finding a route between a source node $s$ and a target node $t$ in a semantic network: (1) semantic navigation (SN), which uses high‐order connectivity via vector embeddings, and (2) shortest‐path (SP) navigation, which finds cost‐minimal routes in the graph.
@@ III.1 Graph representation
The graph is constructed from the initial directed definition links. First, for each headword $i$ defined by a set of $m_{i}$ tokens, a directed edge is drawn from $i$ to each token $j$, with weight $w_{i\to j}=1/m_{i}$. The network is then symmetrized by ensuring that for every edge $i\to j$ there is also a reverse edge $j\to i$. If the reverse edge $j\to i$ is absent, we add it with the same weight $w_{j\to i}:=w_{i\to j}$. If both $i\to j$ and $j\to i$ already exist in the dictionary, we keep both with their original (possibly different) weights, reflecting the distinct outdegrees of $i$ and $j$. This choice is crucial for modeling the real, navigable space of a digital lexicon. For instance, if entry B is in the definition of A, a directed link $A\to B$ exists. However, A may not appear in the definition of B, meaning the reverse link $B\to A$ is absent. A user navigating from A to B can, however, always return to A (e.g., via a “back” button). To capture this fundamental bidirectional navigability, we enforce symmetry. This process reflects the principle that a definitional connection, even if unidirectional, implies a mutual semantic association that can be traversed in both directions. While this symmetric network forms the basis of our main results, we conduct a parallel analysis on the original (non-symmetrized) directed network in the SM [19] to validate our findings under stricter assumptions.
@@ IV.1 Data acquisition and definition processing
Initially, the definitions corresponding to our working lexicon underwent a common set of rigorous filtering and cleaning processes aimed at isolating core semantic relationships and standardizing the lexical base. For entries exhibiting polysemy, we adopted the simplifying convention of using only the first definition provided in the Wiktionary entry. This choice ensures a consistent and reproducible framework for this foundational study. To quantify the impact of this simplification, we conducted a targeted sensitivity analysis (see SM [19], which confirmed that the OD metric is highly sensitive to the specific word sense used. The results show that different senses lead to large, semantically coherent shifts in SOD scores, validating the potential for future work where OD could be extended to handle multiple word senses, for instance by averaging scores or using contextual disambiguation.
Entries’ definitions were then filtered based on their likely part-of-speech (only nouns, verbs, and adjectives retained) to focus on core lexical concepts. Furthermore, heuristics were applied to filter terms potentially representing inflected forms and to identify lemmas for definition tokens, aiming to build the network primarily from base forms. The text was then tokenized into constituent words. These tokens were further filtered: very short or numerical tokens were removed. Finally, to prevent potential trivial results in the subsequent OD calculations, any instance of a headword appearing within its own processed definition was removed.